in order to keep them preserved he plays the heads in an aquarium field with 1000.0 mL of water. If the average human head has a mass of 4500.0 g and a density of 0.6429 g/milliliters what will be the final volume of water after head is placed in the cylinder?
1
Expert's answer
2020-08-25T12:15:18-0400


when a body is immersed in a liquid the volume of the liquid with the body immersed in it increases by the amount of the body volume

"V=\frac{m}{p}=\frac{4500}{0.6429}=6999.53 mL"


V=Vw+Vh=1000+6999.53=7999.53